OpenStudy (anonymous):

please help loves Write the equation of the circle with center (-2, -5) and containing the point (-10, -20)

Directrix (directrix):

To write the equation of the circle, you need the center of the circle, the radius, and the format for writing the circle equation.

Directrix (directrix):

Find the distance between the two points (-2, -5) and (-10, -20) to get the radius. Post that and then we'll write the circle equation. @deszbra

OpenStudy (anonymous):

i used the distance formula and it gave me a number that didnt work

OpenStudy (anonymous):

@Directrix

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

since the center is \((-2,-5)\) you know it has to look like \[(x+2)^2+(y+)^2=r^2\] you just need \(r^2\)

OpenStudy (misty1212):

use the square of the distance formula \[(-2+10)^2+(-5+20)^2\]

OpenStudy (misty1212):

you can almost do it in your head:\[8^2+15^2=64+225=...\]

OpenStudy (anonymous):

thank you